OpenStudy (anonymous):

Please Help me. How to simplify complex fractions? (4+2/xy) /2

sam (.sam.):

\[\large \frac{(\frac{4+2}{xy})}{2}~~~~?\]

OpenStudy (anonymous):

Yes

sam (.sam.):

\[\huge \frac{(\frac{6}{xy})}{2}\] \[\huge \frac{(\frac{6}{xy})\times xy}{2\times xy}\] \[\huge \frac{6}{2xy}\] \[\Huge \frac{3}{xy}\]

OpenStudy (anonymous):

The options were A. xy+1/xy B.2 C.3 D.2xy+1/xy

OpenStudy (anonymous):

Thank You.